A Survey of Study Habits and Attitudes (SSHA) is a psychological test that measures the motivational attitude toward school, and study habits of students. Scores range from 0 to 200. The mean for Canadian college students is about 115, and the standard deviation is about 30. A teacher who suspects this population average to be incorrect gives the SSHA to 25 students who are at least 30 years of age. The sample average score is 127.8. Calculate the p-value.
Answer #1

Note:Final answer is highlighted in colour.

Hypothesis:

H0: \mu_0 =115

Ha: \mu_0 \neq 115

Given,

\overline{X}= 127.8, \sigma = 30, n= 25.

Z= \frac{\overline{X}-\mu_0}{\sigma/\sqrt{n}} = \frac{127.8-115}{30/\sqrt{25}}

= 2.13

p value for Z= 2.13 is 0.033172

We can reject the null hypothesis and conclude that the mean score is different from 115